### 09 Mar 2026 v1

@@ Abstract
Efficient locomotion is important for the evolution of complex life [1, 2], yet the physical principles selecting specific swimming strokes often remain entangled with biological constraints. In viscous fluids, the scallop theorem constrains the temporal organization of strokes [3, 4], but no analogous principle is known for their spatial structure, leaving the prevalence of symmetric gaits across diverse organisms [5, 6] without a physical explanation. Here we show that spatial symmetry acts as an emergent organizing principle for efficiency in viscous fluids. By analysing deformable swimmers whose strokes are not constrained to any particular symmetry class [7, 8], we identify a hydrodynamic duality: symmetric and anti-symmetric strokes are dynamically equivalent, yielding identical speeds and efficiencies, which we prove are optimal among all strokes. By contrast, the optimal efficiency cannot be achieved by generic non-symmetric strokes. We validate this using numerical simulations of Stokes flow [9, 10], demonstrating that these symmetry rules persist even in three-dimensional body plans. Our results suggest that the prevalence of symmetric and alternating gaits in nature reflects not merely a developmental constraint, but a physical optimality principle for locomotion in viscous environments, complementing developmental and neural constraints.
@@ Introduction
Here we address this question by developing a general framework for deformable swimming that treats translation and rotation on equal footing. By relaxing all symmetry constraints, we identify three generic classes of swimming strokes in the case of two dimensions: symmetric, anti-symmetric, and generic non-symmetric (Fig. 1a,b). We show that symmetric and anti-symmetric strokes form a hydrodynamic dual pair: exchanging symmetric deformation modes for their anti-symmetric counterparts leaves the mean swimming speed and efficiency unchanged at leading order, despite distinct surface kinematics. By contrast, the optimal efficiency is always attained by symmetric (or equivalently anti-symmetric) strokes, as generic non-symmetric strokes generate rotational drift that must be actively suppressed at an energetic cost. We validate these findings using numerical simulations of Stokes flow, demonstrating that this symmetry-based organization persists beyond the small-deformation regime and extends to three-dimensional swimmers. Together, our results identify symmetry not as a constraint, but as an emergent organizing principle for efficient locomotion in viscous environments.
@@
A central practical requirement for efficient locomotion is the suppression of rotational drift. While rotation does not contribute directly to net translation, it continuously reorients the swimming direction and therefore reduces the long-time displacement along any fixed axis. Consequently, only straight swimmers with vanishing mean angular velocity ($\langle\Omega\rangle=0$) can achieve high efficiency. The physical mechanism is as follows: generic non-symmetric strokes couple to rotation at $O(\varepsilon^{2})$, the same order as translation. This rotational drift must be suppressed for straight swimming, but doing so requires additional deformation modes that dissipate energy without contributing to propulsion. Rotation thus acts as a parasitic degree of freedom that competes with translation for energetic resources. Importantly, generic non-symmetric deformation protocols generate rotation already at leading order, and straight swimming can only be enforced by introducing additional deformation components that cancel the induced angular velocity. This additional constraint incurs an energetic penalty, so that the optimal efficiency is always attained by symmetry-organized strokes.
Intriguingly, this intuition can be formalized as a rigorous bound. We can prove that among all periodic strokes with vanishing mean rotation, the maximum swimming efficiency is attained by a stroke that is either purely symmetric or purely anti-symmetric. Specifically, for any non-symmetric swimmer tuned to have zero net rotation, one can construct a corresponding pair of symmetric and anti-symmetric strokes that achieves an equal or higher efficiency. Within the space of small-amplitude periodic deformations, the optimal solutions are therefore always contained in the symmetric, and equivalently, anti-symmetric, subspace. We verify this bound numerically by generating random non-symmetric strokes and comparing their efficiencies to those of their symmetry-organized counterparts; as shown in Fig. 2b, symmetric swimmers consistently outperform the original non-symmetric strokes.
Having established that optimal swimming strategies must reside in the symmetric or anti-symmetric subspace, we can then ask what deformation protocol maximizes efficiency. The optimization problem can be solved explicitly; in Fig. 2c, we illustrate the motion of the optimal swimmer when restricting to the second and third, as well as third and fourth Fourier modes. Including higher-order modes further increases the efficiency, which converges to a limiting value $\eta^{*}_{\infty}\approx 0.0563$ (Extended Data Fig. 1). A similar limiting efficiency was first considered by Shapere and Wilczek for symmetric swimmers [8]; we have now proved that there is a corresponding anti-symmetric swimmer that is equally optimal, and that no non-symmetric stroke can surpass this efficiency.
@@ 1.5 Efficiency bound by symmetrized strokes
This proves that the efficiency of an asymmetric stroke cannot exceed that of the more efficient of its two symmetric counterparts. Consequently, the search for an optimal swimming strategy, in terms of efficiency, can be restricted to the subspace of purely symmetric deformations. Note that $\eta\geq\min(\eta^{+},\eta^{-})$ does not generally hold.
@@ 2.6 Stroke protocols and sampling procedures
    Symmetric (n=2,3)(n{=}2,3): $\alpha_{2}(t)=\cos\omega t$, $\beta_{3}(t)=\sin\omega t$.
@@
    Symmetric (n=3,4)(n{=}3,4): $\alpha_{3}(t)=\cos\omega t$, $\beta_{4}(t)=\sin\omega t$.
@@
    Anti-symmetric (n=2,3)(n{=}2,3): $\gamma_{2}(t)=\cos\omega t$, $\delta_{3}(t)=\sin\omega t$.
@@
    Combined: $\alpha_{2}(t)=\cos\omega t$, $\beta_{3}(t)=\sin\omega t$, $\gamma_{2}(t)=\sin\omega t$, $\delta_{3}(t)=\cos\omega t$. This stroke simultaneously activates symmetric and anti-symmetric modes, and also induce rotation.
@@
    Rotational: $\alpha_{3}(t)=\cos\omega t$, $\gamma_{3}(t)=\sin\omega t$.
@@
Duality test with random strokes (Fig. 3b): To test hydrodynamic duality beyond the five specific strokes above, we generated random multi-mode stroke pairs as follows. For each sample, three wavenumbers were drawn uniformly without replacement from $\{2,3,4,5\}$. For each active wavenumber $n$, normal and tangential amplitudes $A_{n}$ and $B_{n}$ were drawn from $\mathrm{Uniform}(-1,1)/(n-1)$ (scaled by 1/(n1)1/(n-1) to moderate the amplitude of higher modes), and phases $\phi^{A}_{n}$, $\phi^{B}_{n}$ were drawn from $\mathrm{Uniform}(0,2\pi)$. The symmetric swimmer was constructed by assigning $\alpha_{n}(t)=A_{n}\cos(\omega t+\phi^{A}_{n})$ and $\beta_{n}(t)=B_{n}\cos(\omega t+\phi^{B}_{n})$; its anti-symmetric dual was obtained via the duality transformation $\alpha_{n}\to-\gamma_{n}$, $\beta_{n}\to\delta_{n}$.
@@
  • Pair 1 (n=2,3n{=}2,3): symmetric $\alpha_{2}(t)=\cos\omega t$, $\beta_{3}(t)=\sin\omega t$; anti-symmetric $\gamma_{2}(t)=\cos\omega t$, $\delta_{3}(t)=\sin\omega t$.
  • Pair 2 (n=3,4n{=}3,4): symmetric $\alpha_{3}(t)=\cos\omega t$, $\beta_{4}(t)=\sin\omega t$; anti-symmetric $\gamma_{3}(t)=\cos\omega t$, $\delta_{4}(t)=\sin\omega t$.

### 10 Mar 2026 v2

@@ Abstract
Efficient locomotion is important for the evolution of complex life [1, 2], yet the physical principles selecting specific swimming strokes often remain entangled with biological constraints. In viscous fluids, the scallop theorem constrains the temporal organization of strokes [3, 4], but no analogous principle is known for their spatial structure, leaving the prevalence of symmetric gaits across diverse organisms [5, 6] without a physical explanation. Here we show that spatial symmetry acts as an emergent organizing principle for efficiency in viscous fluids. By analysing deformable swimmers whose strokes are not constrained to any particular symmetry class [7, 8], we identify a hydrodynamic duality: symmetric and anti-symmetric strokes are dynamically equivalent, yielding identical speeds and efficiencies, which we prove are optimal among all strokes. We validate this using numerical simulations of Stokes flow [9, 10], demonstrating that these symmetry rules persist even in three-dimensional body plans. Our results suggest that the prevalence of symmetric and alternating gaits in nature reflects not merely a developmental constraint, but a physical optimality principle for locomotion in viscous environments, complementing developmental and neural constraints.
@@ Introduction
Here we address this question by developing a general framework for deformable swimming that treats translation and rotation on equal footing. By relaxing all symmetry constraints, we identify three generic classes of swimming strokes in the case of two dimensions: symmetric, anti-symmetric, and generic non-symmetric (Fig. 1a,b). We show that symmetric and anti-symmetric strokes form a hydrodynamic dual pair: exchanging symmetric deformation modes for their anti-symmetric counterparts leaves the mean swimming speed and efficiency unchanged at leading order, despite distinct surface kinematics. By contrast, the optimal efficiency is attained by symmetric (or equivalently anti-symmetric) strokes, as generic non-symmetric strokes generate rotational drift that must be actively suppressed at an energetic cost. We validate these findings using numerical simulations of Stokes flow, demonstrating that this symmetry-based organization persists beyond the small-deformation regime and extends to three-dimensional swimmers. Together, our results identify symmetry not as a constraint, but as an emergent organizing principle for efficient locomotion in viscous environments.
@@
Efficient locomotion also requires suppression of rotational drift: rotation reorients the swimming direction without contributing to net translation, so only straight swimmers with $\langle\Omega\rangle=0$ can achieve high efficiency. Generic non-symmetric strokes couple to rotation at the same order as translation, $O(\varepsilon^{2})$, and suppressing this rotation requires additional deformation modes that dissipate energy without aiding propulsion. Symmetry-organized strokes avoid this competition by construction, making them natural candidates for optimal efficiency.
Intriguingly, this intuition can be formalized as a rigorous bound. We can prove that among all periodic strokes with vanishing mean rotation, no stroke can exceed the maximum swimming efficiency attained within the purely symmetric (or equivalently, purely anti-symmetric) subspace. Specifically, for any non-symmetric swimmer tuned to have zero net rotation, one can construct a corresponding pair of symmetric and anti-symmetric strokes whose efficiencies are at least as high (see Supplementary Information). Remarkably, the bound is tight: there exists a one-parameter family of non-symmetric swimmers that achieves the same efficiency as the symmetric stroke, by combining a symmetric and an anti-symmetric copy of the optimal stroke (see Supplementary Information). Generic non-symmetric strokes, however, do not satisfy this relation; we demonstrate numerically that randomly sampled non-symmetric strokes are consistently outperformed by their symmetry-organized counterparts (Fig. 2b).
Having established that the optimal efficiency can be found within the symmetric (or anti-symmetric) subspace, we can then ask what deformation protocol maximizes efficiency. The optimization problem can be solved explicitly; in Fig. 2c, we illustrate the motion of the optimal swimmer when restricting to the second and third, as well as third and fourth Fourier modes. Including higher-order modes further increases the efficiency, which converges to a limiting value $\eta^{*}_{\infty}\approx 0.0563$ (Extended Data Fig. 1). A similar limiting efficiency was first considered by Shapere and Wilczek for symmetric swimmers [8]; we have now proved that there is a corresponding anti-symmetric swimmer that is equally optimal, and that generic non-symmetric strokes cannot surpass this efficiency.
@@ 1.5 Efficiency bound by symmetrized strokes
This proves that the efficiency of a non-symmetric stroke cannot exceed that of the more efficient of its two symmetric counterparts. Consequently, the search for an optimal swimming strategy, in terms of efficiency, can be restricted to the subspace of purely symmetric deformations. Note that $\eta\geq\min(\eta^{+},\eta^{-})$ does not generally hold.
1.6 Non-symmetric swimmers from matched symmetric–anti-symmetric superposition
Given any symmetric swimmer with deformation amplitudes $(\alpha_{n},\beta_{n})$ and $\gamma_{n}=\delta_{n}=0$, one can construct a one-parameter family of non-symmetric swimmers by adding the anti-symmetric dual stroke with the same shape and phase but scaled by a real parameter $\lambda$:
$\gamma_{m}=\lambda\,\alpha_{m},\qquad\delta_{m}=-\lambda\,\beta_{m}.$ (100)
This superposition preserves the swimming direction: $\langle\Omega\rangle=0$ and $\langle U_{y}\rangle=0$, as can be verified by substituting into Eqs. (73) and (72).
Since the two symmetrized components are $\bm{v}^{\pm}=(1\pm\lambda)\bm{v}$, where $\bm{v}$ is the original symmetric stroke, both yield the same efficiency $\eta^{+}=\eta^{-}=\eta$. It follows that the efficiency bound of the previous subsection is tight: if the original symmetric swimmer achieves $\eta^{\ast}$, the non-symmetric family achieves $\eta^{\ast}$ as well.
Note that this is a specific family of non-symmetric swimmers, which is described by the sum of a symmetric and an anti-symmetric stroke with matched shape and phase. In general, when the symmetric and anti-symmetric components have different shapes or phases, the mismatch generates net rotation ($\langle\Omega\rangle\neq 0$).
@@ 2.6 Stroke protocols and sampling procedures
    Symmetric (n=2,3)(n{=}2,3): $\alpha_{2}(t)=\varepsilon\cos\omega t$, $\beta_{3}(t)=\varepsilon\sin\omega t$.
@@
    Symmetric (n=3,4)(n{=}3,4): $\alpha_{3}(t)=\varepsilon\cos\omega t$, $\beta_{4}(t)=\varepsilon\sin\omega t$.
@@
    Anti-symmetric (n=2,3)(n{=}2,3): $\gamma_{2}(t)=\varepsilon\cos\omega t$, $\delta_{3}(t)=\varepsilon\sin\omega t$.
@@
    Combined: $\alpha_{2}(t)=\varepsilon\cos\omega t$, $\beta_{3}(t)=\varepsilon\sin\omega t$, $\gamma_{2}(t)=\varepsilon\sin\omega t$, $\delta_{3}(t)=\varepsilon\cos\omega t$. This stroke simultaneously activates symmetric and anti-symmetric modes, and also induce rotation.
@@
    Rotational: $\alpha_{3}(t)=\varepsilon\cos\omega t$, $\gamma_{3}(t)=\varepsilon\sin\omega t$.
@@
Duality test with random strokes (Fig. 3b): To test hydrodynamic duality beyond the five specific strokes above, we generated random multi-mode stroke pairs as follows. For each sample, three wavenumbers were drawn uniformly without replacement from $\{2,3,4,5\}$. For each active wavenumber $n$, normal and tangential amplitudes $A_{n}$ and $B_{n}$ were drawn from $\mathrm{Uniform}(-1,1)/(n-1)$ (scaled by 1/(n1)1/(n-1) to moderate the amplitude of higher modes), and phases $\phi^{A}_{n}$, $\phi^{B}_{n}$ were drawn from $\mathrm{Uniform}(0,2\pi)$. The symmetric swimmer was constructed by assigning $\alpha_{n}(t)=\varepsilon A_{n}\cos(\omega t+\phi^{A}_{n})$ and $\beta_{n}(t)=\varepsilon B_{n}\cos(\omega t+\phi^{B}_{n})$; its anti-symmetric dual was obtained via the duality transformation $\alpha_{n}\to-\gamma_{n}$, $\beta_{n}\to\delta_{n}$.
@@
  • Pair 1 (n=2,3n{=}2,3): symmetric $\alpha_{2}(t)=\varepsilon\cos\omega t$, $\beta_{3}(t)=\varepsilon\sin\omega t$; anti-symmetric $\gamma_{2}(t)=\varepsilon\cos\omega t$, $\delta_{3}(t)=\varepsilon\sin\omega t$.
  • Pair 2 (n=3,4n{=}3,4): symmetric $\alpha_{3}(t)=\varepsilon\cos\omega t$, $\beta_{4}(t)=\varepsilon\sin\omega t$; anti-symmetric $\gamma_{3}(t)=\varepsilon\cos\omega t$, $\delta_{4}(t)=\varepsilon\sin\omega t$.